In a square, the sum of the lengths of three sides is 48 and in a rectangle the difference between a length and a width is 7 in. If a width of the rectangle is the same as a side of the square, how long is a length of the rectangle?
Answer by ankor@dixie-net.com(22740)

In a square, the sum of the lengths of three sides is 48 and in a rectangle
the difference between a length and a width is 7 in.
If a width of the rectangle is the same as a side of the square, how long is a length of the rectangle?
:
Let s = side of the square and the width of the rectangle
:
"In a square, the sum of the lengths of three sides is 48"
3s = 48
s = 48/3
s = 16
:
" in a rectangle the difference between a length and a width is 7 in."
L = s + 7
L = 16 + 7
L = 23 is the length of the rectangle